Evaluate the integral of (√siny) cos y dy from 3 to π/2


1
Expert's answer
2021-02-16T12:31:13-0500

Solution.

"\int\limits_3^{\frac{\pi} {2}}\sqrt{sin{y}}\cos{y}dy."


Make a replacement:

"t=\sqrt{\sin{y}}, t\geq0," from here "\sin{y}=t^2."

Then "\cos{y}dy=2tdt."

We will have

"\int\limits_3^{\frac{\pi}{2}}t\cdot 2tdt=\int\limits_3^{\frac{\pi}{2}}2t^2dt=\frac{2t^3}{3}|_3^{\frac{\pi}{2}}=\newline
=\frac{2\sin^{\frac{3}{2}}y}{3}|_3^{\frac{\pi}{2}}=
\frac{2\sin^{\frac{3}{2}}\frac{\pi}{2}}{3}-\frac{2\sin^{\frac{3}{2}}3}{3}=\newline =\frac{2}{3}-\frac{2\sin^{\frac{3}{2}}3}{3}
\approx 0.63."

Answer.

"\int\limits_3^{\frac{\pi} {2}}\sqrt{sin{y}}\cos{y}dy\approx 0.63."